OpenStudy (anonymous):

MEDAL! Yvonne put $4,000 into a savings account. At the end of 3 years, the account had earned $960 simple interest. How many more dollars would she have in her account if the interest rate were 1% greater? Help Please!

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

the interest rate is 8%

OpenStudy (campbell_st):

so 1st you need to find the rate of interest, so use the information you are given formula is I = Prn you know I = 960, P = 4000 and n = 3 so \[960 = 12000r~~~or ~~~r = \frac{960}{12000}\] find r..

OpenStudy (campbell_st):

that will be a decimal value for r, multiply it by 100 to get the percentage value. then add 1 to the answer to find the extra money use the new interest rate as a decimal in the formula with P still 4000 and n still 3 hope it helps

OpenStudy (anonymous):

wait what formula do I use?

OpenStudy (campbell_st):

simple interest I = Prn or I = Prt I = interest, P = Principal, r = interest rate as a decimal n or t = time periods.

OpenStudy (anonymous):

so like this? 960=1200*0.09

OpenStudy (jdoe0001):

well.. we dunno what her interest rate is thus, we'd want to know that first thus \(\bf \qquad \textit{Simple Interest Earned}\\ I = Prt\qquad \begin{cases} I=\textit{interest earned}\to &\$960\\ P=\textit{original amount deposited}\to& \$4000\\ r=rate\\ t=years\to &3 \end{cases} \\ \quad \\ 960=4000\cdot r\cdot 3\impliedby \textit{solve for "r"}\)

OpenStudy (anonymous):

well the thing is I already solved for the interest rate which is 8% and its telling me to add 1% to the interest rate (8%) which all together would be 9%.

OpenStudy (jdoe0001):

right.... so.... you yes.. then you're correct is \(\bf 960=1200*0.09 \) then find the difference between both amounts

OpenStudy (jdoe0001):

hmmm wait a sec....

OpenStudy (jdoe0001):

we know the rate is 8% + 1% = 9% we dunno what the earned interest is so should just be \(\bf i=1200*0.09 \)

OpenStudy (jdoe0001):

well... we're missing he years... so... hmm \(\bf i=1200*0.09 * 3\)

OpenStudy (anonymous):

wait what do we need earned interest for?

OpenStudy (anonymous):

and the years is the 3

OpenStudy (jdoe0001):

well.. at 8%, on a principal of 12,000 over 3 years she's got an earned interest of 960 so...you're asked how much more, would she had now, had it been 9% instead of 8%? so.... find how much is at 8%, we know that, 960 find how much is at 9%, we dunnno that then get their difference and that's how much she would have gottent more, had she been using 9% from the beginning

OpenStudy (anonymous):

I am confused a little, can you set up the equation to plug In the 9% instead of the 8%?

OpenStudy (jdoe0001):

hmmm actually, above I used 12,000, should have been 4000.... but anyway just to clear that out

OpenStudy (anonymous):

ok, can you show me the equation to find how much Yvonne will have in her account if she had 9% as her interest rate?

OpenStudy (jdoe0001):

\(\bf \qquad \textit{Simple Interest Earned}\\ I = Prt\qquad \begin{cases} I=\textit{interest earned}\to &\$960\\ P=\textit{original amount deposited}\to& \$4000\\ r=rate\\ t=years\to &3 \end{cases} \\ \quad \\ 960=4000\cdot r\cdot 3\impliedby r=8\% \\ \quad \\ \quad \\ I = Prt\qquad \begin{cases} I=\textit{interest earned}\\ P=\textit{original amount deposited}\to& \$4000\\ r=rate\to &0.09\\ t=years\to &3 \end{cases} \\ \quad \\ I=4000\cdot 0.09\cdot 3\impliedby r=9\%\)

OpenStudy (anonymous):

I=1080 but that cant be right if her total amount of money in her account is 4,960

OpenStudy (jdoe0001):

1080 is the "earned interest" NOT her account total balance is just the interest earned on the 4000, so the full balance of her account will be 4000+ 1080 at 9% her total account balance will be 4000 + 960 at 8%

OpenStudy (jdoe0001):

if you put $100 in the bank and they pay you 10% after a year that means the earned interest is 10% of 100, or $10 that doesn't mean you put in $100, and after a year you ended up with $10 or $90 less.... simply means, you started with 100, and now added 10 more to it, or 110

OpenStudy (anonymous):

ok now im very confused, can you set up the equation that will show how much Yvonne will have in her account if she had 9% as her interest rate?

OpenStudy (anonymous):

as clear as possible

OpenStudy (jdoe0001):

hmm those are above... can't get clearer I'd think do you even know what an earned interest amount is?

OpenStudy (anonymous):

no, I had not learned about this yet

OpenStudy (jdoe0001):

well... earned interest is amount extra, from the original amount

OpenStudy (anonymous):

oh! wait I do know what it is I just forgot :p

OpenStudy (jdoe0001):

you're asked ok, she made 960 at what rate? you found 8% so at 8% she made 960 at 9% she made how much? what's the difference between the two amounts? or how much would she had made more, had she used 9% instead of 8%

OpenStudy (anonymous):

I know but I don't know how to solve that, that is the reason why I asked this question

OpenStudy (jdoe0001):

well.. how did you get the 8% for the rate? you'd get the "i" for the 2nd rate of 9%, using the same "simple interest formula"

OpenStudy (jdoe0001):

\(\bf \qquad \textit{Simple Interest Earned}\\ I = Prt\qquad \begin{cases} I=\textit{interest earned}\to &\$960\\ P=\textit{original amount deposited}\to& \$4000\\ r=rate\\ t=years\to &3 \end{cases} \\ \quad \\ 960=4000\cdot r\cdot 3\impliedby r=8\% \\ \quad \\ \quad \\ I = Prt\qquad \begin{cases} I=\textit{interest earned}\\ P=\textit{original amount deposited}\to& \$4000\\ r=rate\to &0.09\\ t=years\to &3 \end{cases} \\ \quad \\ I=4000\cdot 0.09\cdot 3\impliedby r=9\%\)

OpenStudy (anonymous):

I got 8% for the rate by using the formula to find the rate --- A=1/P*T *100

OpenStudy (jdoe0001):

well... you got 1080 already from the 9%, right? so.. we know at 8% she earned 960 we also know at 9% she earned 1080 so 1080 - 960 = " how much MORE she would have gotten, had she used 9%"

OpenStudy (anonymous):

ok now this makes much more sense